OpenStudy (anonymous):

(x2 + 4x − 3)(2x2 + x + 6). can i get help plz? i will fan and medal i do not under stand this

OpenStudy (anonymous):

choices Find the product of (x2 + 4x − 3)(2x2 + x + 6). 2x4 + 9x3 + 4x2 + 21x − 18 2x4 + 9x3 + 10x2 + 21x − 18 2x4 + 7x3 + 4x2 + 21x − 18 2x4 + 9x3 + 4x2 + 27x − 18

OpenStudy (anonymous):

i got 2x^4+4x^2-18

OpenStudy (anonymous):

Ready?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

(x2+4x−3)(2x2+x+6) =(x2+4x+−3)(2x2+x+6) =(x2)(2x2)+(x2)(x)+(x2)(6)+(4x)(2x2)+(4x)(x)+(4x)(6)+(−3)(2x2)+(−3)(x)+(−3)(6) =2x4+x3+6x2+8x3+4x2+24x−6x2−3x−18 =2x4+9x3+4x2+21x−18

OpenStudy (anonymous):

so is the answer this 2x4+9x3+4x2+21x18

OpenStudy (anonymous):

so a right ?

OpenStudy (anonymous):

Yea. If I didn't make a mistake on the paper. I hope i got it right. Sorry if i get it wrong

OpenStudy (anonymous):

thx it is ok